Annuity A makes annual payments of $813.73 for each of the next 10 years, while annuity B makes annual payments of $500 per year forever. At what interest rate would you be indifferent between the two? At interest rates above/below this break-even rate, which annuity would you choose?
Answer #1

813.73/r[1 - (1 + r)-10] = 500/r

[1 - (1 + r)-10] = 0.6145

1 - 0.6145 = (1 + r)-10

(1 + r)10 = 2.594

r = 10.00%